\begin{thebibliography}{99}

\bibitem{barros2017first}
Barros, L. C., Bassanezi, R. C., Lodwick, W. A.  
\textit{A first course in fuzzy logic, fuzzy dynamical systems, and biomathematics: theory and applications.}  
Springer, 2017.

\bibitem{de2015diferenccas}
Barros, L. C., Santo Pedro, F., Gomes, L. T.  
"Diferenças entre números fuzzy."  
\textit{Proceeding Series of the Brazilian Society of Computational and Applied Mathematics}, vol. 3, no. 1, 2015.

\bibitem{DEBARROS201764}
Barros, L. C., Santo Pedro, F.  
"Fuzzy differential equations with interactive derivative."  
\textit{Fuzzy Sets and Systems}, vol. 309, pp. 64--80, 2017.  
doi:10.1016/j.fss.2016.04.002.


\bibitem{fcorrelacionado}
Cabral, V. M., Prata, R. A. C., Barros, L. C.  
"Operações com números fuzzy f-correlacionados."  
\textit{Biomatemática}, vol. 23, pp. 15--32, 2013.

\bibitem{1375791}
Carlsson, C., Fullér, R., Majlender, P.  
"Additions of completely correlated fuzzy numbers."  
\textit{2004 IEEE International Conference on Fuzzy Systems (IEEE Cat. No.04CH37542)}, vol. 1, pp. 535--539, 2004.  
doi:10.1109/FUZZY.2004.1375791.

\bibitem{carlsson2004additions}
Carlsson, C., Fullér, R., others.  
"Additions of completely correlated fuzzy numbers."  
\textit{2004 IEEE International Conference on Fuzzy Systems (IEEE Cat. No. 04CH37542)}, vol. 1, pp. 535--539, 2004.

\bibitem{coroianu2013additivity}
Coroianu, L., Fullér, R.  
"On additivity of the weighted possibilistic mean operator."  
\textit{2013 IEEE 14th International Symposium on Computational Intelligence and Informatics (CINTI)}, pp. 303--308, 2013.

\bibitem{coroianu2013multiplication}
Coroianu, L., Fullér, R.  
"On multiplication of interactive fuzzy numbers."  
\textit{2013 IEEE 11th International Symposium on Intelligent Systems and Informatics (SISY)}, pp. 181--185, 2013.

\bibitem{fuller2004interactive}
Fullér, R., Majlender, P.  
"On interactive fuzzy numbers."  
\textit{Fuzzy Sets and Systems}, vol. 143, no. 3, pp. 355--369, 2004.

\bibitem{FULLER2003363}
Fullér, R., Majlender, P.  
"On weighted possibilistic mean and variance of fuzzy numbers."  
\textit{Fuzzy Sets and Systems}, vol. 136, no. 3, pp. 363--374, 2003.  
doi:10.1016/S0165-0114(02)00216-6.

\bibitem{FULLER201150}
Fullér, R., Mezei, J., Várlaki, P.  
"An improved index of interactivity for fuzzy numbers."  
\textit{Fuzzy Sets and Systems}, vol. 165, no. 1, pp. 50--60, 2011.  
doi:10.1016/j.fss.2010.06.001.


\bibitem{ZADEH1965338}
Zadeh, L. A.  
"Fuzzy sets."  
\textit{Information and Control}, vol. 8, no. 3, pp. 338--353, 1965.  
doi:10.1016/S0019-9958(65)90241-X.

\bibitem{ZADEH1975199}
Zadeh, L. A.  
"The concept of a linguistic variable and its application to approximate reasoning—I."  
\textit{Information Sciences}, vol. 8, no. 3, pp. 199--249, 1975.  
doi:10.1016/0020-0255(75)90036-5.

\end{thebibliography}

\documentclass[a4paper, 10pt, oneside]{article}
\usepackage{times}
\usepackage{amsfonts,amsmath,amssymb,amsthm}
\usepackage{graphicx} 
\newtheorem{df}{Definition}
\newtheorem{teo}{Theorem}
\newtheorem{ex}{Example}
\newtheorem{lem}{Lemma}
\newtheorem{cor}{Corollary}

\title{Properties of $f$ correlated fuzzy numbers}
\author{Diogo Sampaio da Silva \& Roberto Antonio Cordeiro Prata }
\date{\today}

\begin{document}

\maketitle

\begin{abstract}
This paper presents some concepts of the theory of interactive fuzzy numbers, and mainly, a class of interactive fuzzy numbers, called $f$-correlated fuzzy numbers. We start from the foundations of general fuzzy mathematics and go through operations and the notion of interactivity for fuzzy numbers. The main result is that $f$-correlation preserve the shape of certains fuzzy numbers. More specificaly, if two fuzzy numbers are $f$ correlated, and one is a LR-type fuzzy number, the other is also a LR-type fuzzy number. This paper also presents some operations with the $f$-correlated fuzzy numbers wich are interesting to applications like biomathematics.
\end{abstract}



\section{Introduction}
A fuzzy set, as opposed to a classical or crisp set, is a set with continuous degrees of membership, modeling real situations where there is no exact criterion delineating a class of objects \cite{barros2017first}. Therefore, a fuzzy set is represented by a function in the form $\varphi_{A} : U \to [0,1]$, while a classic set is represented by its characteristic function, i.e., a function $\chi_{B} :U \to \left\{ 0,1\right\}$ such that $B = \left\{ x \in U; \chi_{B}(x) =1 \right\}$, being a particular case of a fuzzy set. Here, the domain is a universal set and the function's output represents the degree of membership, increasing progressively.

In a similar way, the notion of fuzzy number was developed to generalize the concept of a real number. By applying Zadeh's Extension Principle, we derive operations among fuzzy numbers, utilizing closed intervals of real numbers \cite{barros2017first}. All these ideas come from the origins of Fuzzy Mathematics, in the second half of the 20th century, from the work of Zadeh \cite{ZADEH1965338}. The initial idea was to expand the notion of set, traditionally defined by the dichotomy between membership and non-membership. Thus, to formally deal with propositions and uncertain quantities, the concept of fuzzy set was introduced along with some operations of fuzzy sets, their properties and also how to define functions with fuzzy sets. Later, Zadeh introduced the notion of non-interactivity to deal with approximate reasoning \cite{ZADEH1975199}. We use this notion to  study a class of fuzzy numbers, called interactive fuzzy numbers. And among the interactive fuzzy numbers, the $f$-correlated fuzzy numbers are our object of interest.

To define fuzzy numbers we must define $\alpha$-levels of fuzzy sets. If the universe $U$ is a topological space, then the $\alpha$-level of $A$ is $[A]^{\alpha} = \left\{ a \in U; \varphi_{A} \ge \alpha \right\}$. The support of $A$ is $\text{supp} A = \left\{ a \in U; \varphi_{A} > 0 \right\}$. And for $\alpha = 0$, the $\alpha$-level of $A$ is $[A]^{\alpha} = \overline{\text{supp} A}$ \cite{barros2017first}. An useful fact about $\alpha$-levels is that they determine fuzzy sets uniquely.

\begin{lem}[\cite{barros2017first}]
Let \( A \) and \( B \) be fuzzy subsets of \( U \). A necessary and sufficient condition for \( A = B \) to hold is that \( [A]_\alpha = [B]_\alpha \) for all \( \alpha \in [0, 1] \).
\end{lem}

\begin{proof}
Naturally \( A = B \implies [A]_\alpha = [B]_\alpha \) for all \( \alpha \in [0, 1] \). Now, we suppose that \( [A]_\alpha = [B]_\alpha \) for all \( \alpha \in [0, 1] \). If \( A \neq B \), then there is an \( x \in U \) such that \( \varphi_A(x) \neq \varphi_B(x) \). Therefore, we have \( \varphi_A(x) < \varphi_B(x) \) or, conversely, \( \varphi_A(x) > \varphi_B(x) \).

If we assume that \( \varphi_A(x) > \varphi_B(x) \), then we come to the conclusion that \( x \in [A]_{\varphi_A(x)} \) and \( x \notin [B]_{\varphi_A(x)} \), and therefore \( [A]_{\varphi_A(x)} \neq [B]_{\varphi_A(x)} \), which contradicts the hypothesis that \( [A]_\alpha = [B]_\alpha \) for all \( \alpha \in [0, 1] \). A similar contradiction is reached if we assume that \( \varphi_A(x) < \varphi_B(x) \).
\end{proof}

\begin{df}[Fuzzy number \cite{barros2017first}]
A fuzzy subset $A$ of $\mathbb{R}$ is a fuzzy number if every $\alpha$-level is a non-empty bounded closed interval of $\mathbb{R}$ with closure of the support.
\end{df}

\begin{df}[LR-type fuzzy number \cite{FULLER2003363}]
A LR-type fuzzy number \( A \) is identified with the following membership function:

\[
A(u) = 
\begin{cases}
L\left( \dfrac{q_{-} - u}{a} \right) & \text{if } u \in [q_{-} - a , q_{-}], \\
1 & \text{if } u \in [q_{-}, q_{+}], \\
R\left( \dfrac{u - q_{+}}{b} \right) & \text{if } u \in [q_{+} , q_{+} + b], \\
0 & \text{otherwise},
\end{cases}
\]
where \( [q_{-}, q_{+}] \) is the peak of fuzzy number \( A \); \( q_{-} \) and \( q_{+} \) are the lower and upper modal values; \( L, R : [0, 1] \rightarrow [0, 1] \) with \( L(0) = R(0) = 1 \) and \( L(1) = R(1) = 0 \) are non-increasing, continuous functions. We will use the notation \( A = (q_{-}, q_{+}, a, b)_{LR} \).
\end{df}

Hence, the closure of the support of \( A \) is exactly \( [q_{-} - a, q_{+} + \beta] \). If \( L \) and \( R \) are strictly decreasing functions, then the \( \alpha \)-level sets of \( A \) can easily be computed as $[A]^\alpha = [q_{-} - a L^{-1}(\alpha), q_{+} + b R^{-1}(\alpha)], \alpha \in [0, 1]$.

\begin{ex}[Trapezoidal and triangular fuzzy numbers]
A Trapezoidal fuzzy number $T$ is defined as a LR-Type fuzzy number with functions $L\left( \dfrac{q_{-} - u}{a} \right) = \frac{u+a-q_{-}}{a}$ and $R\left( \dfrac{u - q_{+}}{b} \right) = \frac{q_{+} + b - u}{q_{+} -u}$, i.e., $T$ is identified by the following membership function:

\[
T(u) = 
\begin{cases}
\frac{u+a-q_{-}}{a} & \text{if } u \in [q_{-} - a , q_{-}], \\
1 & \text{if } u \in [q_{-}, q_{+}], \\
\frac{q_{+} + b - u}{q_{+} -u} & \text{if } u \in [q_{+} , q_{+} + b], \\
0 & \text{otherwise},
\end{cases}
\]
where \( [q_{-}, q_{+}] \) is the peak of fuzzy number \( A \); \( q_{-} \) and \( q_{+} \) are the lower and upper modal values. If $q_{-} = q_{+}$, then $T$ is a triangular fuzzy number.
\end{ex}

Through the joint possibility distribution of two fuzzy numbers, we explore the so-called interactive fuzzy numbers, which can be described as those whose values are interdependently assigned \cite{fuller2004interactive}.

\begin{df}[Possibility Distribution \cite{de2015diferenccas}]
A possibility distribution over $\Omega \neq \varnothing$ is a function $\varphi : \Omega \rightarrow [0,1]$ that satisfies $\sup_{\omega \in \Omega} \varphi (\omega) = 1$.
\end{df}

\begin{df}[Joint Possibility Distribution \cite{FULLER201150}]
Let $A$ and $B$ be fuzzy numbers and $C \in F_{C}(\mathbb{R}^2)$. Then, $\varphi_{C}$ is a joint possibility distribution of $A$ and $B$ if:
$$\max_{y \in \mathbb{R}} \varphi_{C}(x,y) = \varphi_{A}(x) \text{ and } \max_{x \in \mathbb{R}} \varphi_{C}(x,y) = \varphi_{B}(y).$$
Moreover, $\varphi_{A}$ and $\varphi_{B}$ are called the marginal distributions of $C$.
\end{df}

\begin{df}[Extension principle with joint possibility distributions \cite{1375791}]
Let $C$ be the joint possibility distribution of (marginal possibility distributions) $A_1, \ldots, A_n \in \mathcal{F}$, and let $f : \mathbb{R}^n \rightarrow \mathbb{R}$ be a continuous function. Then $f_C(A_1, \ldots, A_n)$ will be defined by

\[
f_C(A_1, \ldots, A_n)(y) = \sup_{y = f(x_1, \ldots, x_n)} C(x_1, \ldots, x_n). \quad (1)
\]
\end{df}

\begin{lem}[\cite{1375791}]
Let \( A_1, \ldots, A_n \) be fuzzy numbers, let \( C \) be their joint possibility distribution, and let \( f : \mathbb{R}^n \to \mathbb{R} \) be a continuous function. Then,

\[
\left [ f_{C} (A_1, \ldots, A_n) \right ]^{\alpha} = f([C]^{\alpha})
\]

for all \( \alpha \in [0, 1] \).
\end{lem}

\begin{proof}
For any \( \alpha \in [0, 1] \) we obtain:
\begin{align*}
\left [ f_{C} (A_1, \ldots, A_n) \right ]^{\alpha} &= \overline{\left\{ y \in \mathbb{R} ; \sup_{y=f(x)}  C(x) > \alpha \right\}}\\ 
 &= \overline{\left\{ y \in \mathbb{R} ; \exists x \in \mathbb{R} ; y = f(x), C(x) > \alpha \right\}}\\ 
 &= \overline{\left\{ f(x) \in \mathbb{R} ; C(x) > \alpha \right\}}\\ 
 &= f \left ( \overline{\left\{ x \in \mathbb{R}^{n} ; C(x) > \alpha \right\}} \right ) \\
 &= f([C]^{\alpha })
\end{align*}
This concludes our proof.
\end{proof}

\begin{df}
Fuzzy numbers $A$ and $B$ are said to be non-interactive if and only if their joint possibility distribution $C$ satisfies the relation $$\forall \, x,y \in \mathbb{R}, \varphi_{C} (x,y) = \min \left \{ \varphi_{A} (x), \varphi_{B} (y) \right \}.$$ Otherwise, they are said to be interactive.
\end{df}

If $A, B \in F(\mathbb{R})$ are non-interactive, then the joint membership function is given by $A \times B$. In this case, changes in the membership function of $A$ will not alter the second marginal possibility distribution and vice versa. Thus, we have the equality $[B]^{\alpha}=[A_{1}]^{\alpha} \times  [A_{2}]^{\alpha}, \forall \, \alpha \in [0,1]$. Consequently, $A, B \in F(\mathbb{R})$ are said to be interactive if they cannot take their values independently of each other.

If $A_1, \ldots, A_n$ are non-interactive, i.e., their joint possibility distribution is given by $C(x_1, \ldots, x_n) = \min \{A_1(x_1), \ldots, A_n(x_n)$ then we obtain the Zadeh's extension principle, and if $C(x_1, \ldots, x_n) = T(A_1(x_1), \ldots, A_n(x_n))$ where $T$ is a t-norm, then we get the t-norm-based extension principle \cite{1375791}.

The concept of $f$-correlated fuzzy numbers is a generalization of the concept of linearly correlated fuzzy numbers, now using a monotonic injective function instead of a linear function. An interesting particular case is when the correlation function $f$ is hyperbolic \cite{fcorrelacionado}.

\begin{df}[$f$-Correlated Fuzzy Numbers \cite{fcorrelacionado}]
Let $X,Y \subset \mathbb{R}$ and $f: X \to Y$ be a monotonic injective and continuous function. Two fuzzy numbers $A$ and $B$ are correlated according to the function $f$ or $f$-correlated if their joint possibility distribution $J$ is given by

\begin{align*}
\varphi_{J}(x,y) &= \varphi_{A}(x) \chi_{y = f(x)} (x,y)\\ 
 &= \varphi_{B}(y) \chi_{y = f(x)} (x,y),
\end{align*}
where $\chi_{y = f(x)} (x,y)$ is the characteristic function of the graph of $f$, or, $\left \{ (x,y) \in \mathbb{R}^{2}; y = f(x) \right \}$. 
\end{df}

\begin{ex}[Completely Correlated Fuzzy Numbers \cite{1375791}]
Two fuzzy numbers $A$ and $B$ are declared completely correlated if there exists $q, r \in \mathbb{R}$, with $q \neq 0$, such that their joint possibility distribution, with $\chi_{qx+r=y}(x,y)$ being the characteristic function of the line $\left \{ (x, y) \in \mathbb{R}^2 ; qx + r = y \right \}$, is defined by:
$$\varphi_{C}(x, y) = \varphi_{A}(x) \chi_{qx+r=y} (x, y) = \varphi_{B}(y) \chi_{qx+r=y} (x, y).$$
\end{ex}

\begin{ex}[Hyperbolically Interactive Fuzzy Numbers]
Consider $f$ being a hyperbolic function $f: \mathbb{R} \longrightarrow \mathbb{R}$, given by $f(x) = \dfrac{q}{x} + r, x \ne 0$. Two fuzzy numbers $A$ and $B$ are hyperbolically interactive if there exist $q,r \in \mathbb{R}$, $q \ne 0$, such that their joint possibility distribution $J$ is given by

\begin{align*}
\varphi_{J}(x,y) &= \varphi_{A}(x) \chi_{\dfrac{q}{x} + r = y} (x,y)\\ 
 &= \varphi_{B}(y) \chi_{\dfrac{q}{x} + r = y} (x,y),
\end{align*}
where $\chi_{\dfrac{q}{x} + r = y} (x,y)$ is the characteristic function of the set $\left \{ (x,y) \in \mathbb{R}^{2}; \dfrac{q}{x} + r =y \right \}$. In this case, $[B]^{\alpha}=\left \{ \dfrac{q}{x} + r; x \in [A]^{\alpha} \right \}$, $0 \notin [A]^0$.
\end{ex}

\section{Results and Discussions}
\subsection{Main results}
It is important to remebember that continuous monotone injective functions have the following property: $f([a,b]) = [f(a), f(b)]$, if $f$ is increasing, and $f([a,b]) = [f(b), f(a)]$ , if $f$ is decreasing.

\begin{teo}
If $A$ is a $LR$-type fuzzy number, with $L$ and $R$ injective functions, and $B$ is $f$-correlated to $A$, then for all $\alpha \in [0,1]$, the extremes of the $[B]^{\alpha}$ are the result of applying $f$ in the extremes of $[A]^{\alpha}$.
\end{teo}

\begin{proof}
Since for all $\alpha \in [0,1]$, $[B]^{\alpha} = f([A]^{\alpha}) = f([q_{-} - a L^{-1}(\alpha), q_{+} + b R^{-1}(\alpha)])$, then if $f$ is increasing,  $[B]^{\alpha} = \left[f\left(q_{-} - a L^{-1}(\alpha)\right), f \left( q_{+} + b R^{-1}(\alpha) \right) \right]$ and if $f$ is decreasing, $[B]^{\alpha} = \left[f \left( q_{+} + b R^{-1}(\alpha) \right) , f\left(q_{-} - a L^{-1}(\alpha)\right) \right]$.
\end{proof}

\begin{cor}
If $A$ is a $LR$-type fuzzy number, with $L$ and $R$ injective functions, and $B$ is completly correlated to $A$, then $B$ is also a $LR$-type fuzzy number.
\end{cor}
\subsection{Examples}
We start with some observations about properties of classes of $f$ correlated fuzzy numbers. A interesting property of completely correlated fuzzy numbers is that they preserve the shape of each other, in the sense we now investigate.
\begin{ex}
If $A$ is is a triangular fuzzy number, completely correlated with a fuzzy number $B$, then $B$ is also a triangular fuzzy number. In fact, for \(\alpha \in [0, 1]\), the alpha-level of \(A\) is $[A]^\alpha = [a + \alpha (u - a),  b - \alpha (u - b)]$, where $(a,0)$, $(b,0)$ and $(u,1)$ are the vertices of a triangle in the Cartesian plane. Since $[B]^\alpha = f([A]^\alpha)$, we deduce $[B]^\alpha = f[a + \alpha (u - a),  b - \alpha (u - b)] = [f(a + \alpha (u - a)), f(b - \alpha (u - b))] = [(qa+r) + \alpha (qu+r - qa-r),  (qb+r) - \alpha (qu+r - qb-r)]$, which is exactly the $\alpha$-level of the triangular fuzzy number with vertices $(qa+r,0)$, $(qb+r,0)$ and $(qu+r,1)$.
\end{ex}

\begin{ex}
If $A$ is is a trapezoidal fuzzy number, completely correlated with a fuzzy number $B$, then $B$ is also a trapezoidal fuzzy number.  Indeed, for \(\alpha \in [0, 1]\), the alpha level of \(A\) is $[A]^\alpha = [ a + \alpha (b - a), d + \alpha (c - d)]$, where $(a,0)$, $(b,1)$, $(c,1)$ and $(d,0)$ are the vertices of a trapezoid in the Cartesian plane. Because $[B]^\alpha = f([A]^\alpha)$, we obtain $[B]^\alpha = f[ a + \alpha (b - a), d + \alpha (c - d)] = [ f(a + \alpha (b - a)), f(d + \alpha (c - d))] = [ (qa+r) + \alpha (qb+r - qa-r) , (qd+r) - \alpha (qc+r - qd-r)]$, which is exactly the $\alpha$-level of the trapezoidal fuzzy number with vertices $(qa+r,0)$, $(qb+r,1)$, $(qc+r,1)$ and $(qd+r,0)$.
\end{ex}

\end{document}